OpenStudy (anonymous):

Please Help! A sample of helium gas has a volume of 620 ml at a temperature of 500k. If we decrease the temperature to 100K while keeping the pressure constant what will the new volume be?

OpenStudy (anonymous):

\[At constant temperature, we can use the relation ( assuming the gas is an ideal gas), V1/T1 = V2/T2, where V1= initial volume T1 = initial temperature V2 = final volume T2 = final temperature For temperature should be in kelvin.

OpenStudy (anonymous):

so you would cross multiply

OpenStudy (anonymous):

Ya you have to cross multiply. By the way, its at constant pressure and not at the constant temperature the relation is valid. In your question, The initial volume is 620ml and initial temperature is 500k. Then the final temperature is 100K. So, the new volume will be V2 given by:V2 = \[\frac{ V1 }{ T1 }XT2\] Now all you have to do is substitute the values. But the best way to understand these type of questions is deriving the ideal gas equation\[PV = nRT\] and derive the different relations at constant temperature, volume and pressure.

OpenStudy (anonymous):

\[\frac{ 620ml }{500k }=\frac{ ? }{ 100k}\]

OpenStudy (anonymous):

620x100=62000 62000/500 =124ML

OpenStudy (anonymous):

Ya, thats pretty much it.